### v2

@@ 2.4 Error Propagation in Integration
$\Phi_{f}(Z_{n}+E_{n})\approx\Phi_{f}(Z_{n})+\frac{\partial\Phi_{f}}{\partial Z}\bigg{|}_{Z_{n}}E_{n}\;,$
where $\tfrac{\partial\Phi_{f}}{\partial Z}\big{|}_{Z_{n}}$ denotes the Jacobian of the learned flow map $\Phi_{f}$ with respect to the state, evaluated at $Z_{n}$. Inserting this into the expression for $E_{n+1}$ gives
$E_{n+1}\approx\underbrace{\big{(}\Phi_{f}(Z_{n})-\Phi_{g}(Z_{n})\big{)}}_{\text{local error}}+\underbrace{\frac{\partial\Phi_{f}}{\partial Z}\bigg{|}_{Z_{n}}E_{n}}_{\text{propagated error}}.$
Thus, each step combines a local error (discrepancy between $\Phi_{f}$ and $\Phi_{g}$) and a propagated error (amplification of the previous error by the solver Jacobian), showing how small inaccuracies in the learned acceleration field can accumulate over time and underscoring the need for accurate acceleration targets during training.
@@ 2.5 Loss Function and Optimization
The FNODE loss directly supervises accelerations. Given predicted accelerations $\hat{\ddot{\mathbf{z}}}_{i}=f(\mathbf{z}_{i},\dot{\mathbf{z}}_{i};\mathbf{\Theta})$ and reference accelerations $\ddot{\mathbf{z}}_{i}$ from trajectory data, we minimize the mean squared error (MSE)
$\mathcal{L}(\mathbf{\Theta})=\frac{1}{N}\sum_{i=0}^{N-1}\lVert\hat{\ddot{\mathbf{z}}}_{i}-\ddot{\mathbf{z}}_{i}\rVert^{2}_{2}.$ (24a)
@@
$\mathbf{\Theta}^{*}=\arg\min_{\mathbf{\Theta}}\mathcal{L}(\mathbf{\Theta}).$ (24b)
We train FNODE using the Adam optimizer, which provides stable and efficient convergence in practice. The training algorithm is shown in Appendix A. A discussion of the training costs is deferred to Section 4.
@@ 3.1 Single-Mass-Spring-Damper System
As a first benchmark, we consider a simple single-mass-spring-damper system. Its dynamics are governed by
@@
with $m=10$ kg, $k=50$ N/m, and $d=2$ Ns/m. The system is initialized with $x(0)=1$ m, $\dot{x}(0)=0$, and simulated with RK4 at $\Delta t=0.01$ s. A trajectory of 700 steps was used for training, with an additional 300 steps reserved for extrapolation. Model hyperparameters are listed in Table 2.
@@
Figure 2 compares predicted displacement and velocity. All methods fit well within the training interval, but differences emerge in extrapolation. FNODE maintains close agreement with the ground truth ($\epsilon=1.9\text{e-}2$), while MBD-NODE degrades more quickly ($\epsilon=1.3\text{e-}1$). LSTM fails to capture the dissipative behavior ($\epsilon=7.0\text{e-}1$), and FCNN diverges entirely ($\epsilon=4.2\text{e}1$). Phase-space trajectories in Fig. 3 confirm FNODE’s stable long-term prediction, in contrast to the drift or stagnation observed in the baselines.
Figure 2: Temporal evolution of state variables for the single-mass-spring-damper system. The left and right columns depict the displacement ($x$) and velocity ($v$), respectively. Dashed lines indicate model predictions over the training interval ($t\in[0,3]$), while dotted lines represent extrapolated predictions on the test interval. The mean squared error (MSE) for each model is provided: FNODE ($\epsilon=1.9\text{e-}2$), MBD-NODE ($\epsilon=1.3\text{e-}1$), LSTM ($\epsilon=7.0\text{e-}1$), and FCNN ($\epsilon=4.2\text{e}1$).
Figure 3: Comparative phase-space trajectories ($v$ vs. $x$) for the single-mass-spring-damper system. Model performance on the training and test data is differentiated by dashed and dotted lines, respectively. The subplots correspond to the predictions of (a) FNODE, (b) MBD-NODE, (c) LSTM, and (d) FCNN.
@@ 3.2 Multiscale Triple-Mass-Spring-Damper System
3.3 Triple-Mass-Spring-Damper System
We next consider a three-mass-spring-damper system with strong scale separation (Fig. 4). The masses are $m_{1}=100$ kg, $m_{2}=10$ kg, and $m_{3}=1$ kg. All springs have stiffness $k_{i}=50$ N/m and all dampers $d_{i}=2$ Ns/m. The governing equations are
$\ddot{x}_{1}$ $=-\tfrac{k_{1}}{m_{1}}x_{1}-\tfrac{d_{1}}{m_{1}}(\dot{x}_{1}-\dot{x}_{2})+\tfrac{k_{2}}{m_{1}}(x_{2}-x_{1})+\tfrac{d_{2}}{m_{1}}(\dot{x}_{2}-\dot{x}_{1}),$ (26)
$\ddot{x}_{2}$ $=-\tfrac{k_{2}}{m_{2}}(x_{2}-x_{1})-\tfrac{d_{2}}{m_{2}}(\dot{x}_{2}-\dot{x}_{1})+\tfrac{k_{3}}{m_{2}}(x_{3}-x_{2})+\tfrac{d_{3}}{m_{2}}(\dot{x}_{3}-\dot{x}_{2}),$
$\ddot{x}_{3}$ $=-\tfrac{k_{3}}{m_{3}}(x_{3}-x_{2})-\tfrac{d_{3}}{m_{3}}(\dot{x}_{3}-\dot{x}_{2}).$
Figure 4: Triple mass-spring-damper system. The setup is similar to the single mass-spring-damper system, except for the addition of two more masses, springs, and dampers.
The system was simulated with RK4 at a fixed step of $\Delta t=0.01$ s. Training trajectories consisted of 300 steps from the initial state $x_{1}=1$, $x_{2}=2$, $x_{3}=3$, $v_{1}=v_{2}=v_{3}=0$ (SI units). Performance was then evaluated by extrapolation for an additional 100 steps. Model hyperparameters are listed in Table 3.
@@ 3.3 Triple-Mass-Spring-Damper System
Figure 5 shows the trajectories of $x_{1},x_{2},x_{3}$ and $v_{1},v_{2},v_{3}$ during training and testing. All models fit the training data, though FCNN exhibits minor oscillations. In extrapolation ($t>3$), FNODE achieves the best agreement with ground truth ($\epsilon=3.0\text{e-}2$). MBD-NODE remains reasonable but drifts due to solver error accumulation. LSTM largely repeats training patterns and fails to generalize, while FCNN diverges under multiscale dynamics.
Figure 6 shows phase-space trajectories on the test set. FNODE tracks the dynamics well, though small discrepancies appear for the lighter masses due to irregular acceleration fields and error accumulation. MBD-NODE performs poorly for Mass 1, where low variance in the signal reduces its contribution to the loss. LSTM largely repeats training patterns and fails to generalize, while FCNN is inaccurate both in- and out-of-distribution.
Figure 5: Predicted temporal evolution for the triple-mass-spring-damper benchmark. The left column presents the displacement ($x$) and the right column presents the velocity ($v$) for each of the three masses. Dashed and dotted lines denote predictions on the training ($t\in[0,3]$) and test sets, respectively. The associated MSEs are: FNODE ($\epsilon=3.0\text{e-}2$), MBD-NODE ($\epsilon=5.0\text{e-}2$), LSTM ($\epsilon=1.0\text{e}0$), and FCNN ($\epsilon=1.3\text{e}0$).
@@ 3.4 Double Pendulum
We next evaluate FNODE on a chaotic benchmark: the double pendulum with $m_{1}=m_{2}=1$ kg, $l_{1}=l_{2}=1$ m, and $g=9.81$ m/s2. The dynamics follow from the Hamiltonian formulation (see Appendix B for details).
The double pendulum was simulated with $L_{1}=L_{2}=1$ m, $m_{1}=m_{2}=1$ kg, and $g=9.81$ m/s2. The initial state was $\theta_{1}=3\pi/7$, $\theta_{2}=3\pi/4$, and $\dot{\theta}_{1}=\dot{\theta}_{2}=0$. Trajectories were generated using RK4 with $\Delta t=0.01$ s for 300 training steps and 100 extrapolation steps. Model hyperparameters are listed in Table 4.
@@
Max. epochs 10000 450 400 600
Initial learning rate 1e-3 1e-3 5e-4 5e-4
Learning rate decay 0.7 0.98 0.98 0.99
@@
Figure 8 shows the double pendulum trajectories. All models fit the training data, but diverge in extrapolation. FNODE, despite error accumulation in the ODE solver, achieves the best accuracy ($\epsilon=1.9\text{e-}2$).
Phase-space plots in Fig. 9 confirm this trend: FNODE reproduces the main chaotic patterns, with only minor discrepancies. MBD-NODE loses accuracy in testing, LSTM largely repeats training patterns without generalization, and FCNN fails to capture the dynamics outside the training regime.
Figure 8: Dynamic response predictions for the double pendulum system. The left and right columns show the temporal evolution of angular position ($\theta$) and angular velocity ($\omega$), respectively. Dashed lines represent the fit to the training data ($t\in[0,3]$), while dotted lines show the extrapolation on test data. The reported MSE values are: FNODE ($\epsilon=1.9\text{e-}2$), MBD-NODE ($\epsilon=3.7\text{e-}1$), LSTM ($\epsilon=2.2\text{e}0$), and FCNN ($\epsilon=8.7\text{e}0$).
@@
3.5 Slider Crank
Figure 10: Slider-crank mechanism with motor torque $\tau_{1}$, rotational dampers at all joints, and slider friction.
We next test long-term prediction on a slider–crank mechanism (Fig. 10). The system is simulated for 45 s (4500 steps), with 1500 steps used for training and the remainder for testing. The mechanism consists of three rigid bodies with a slider connected to a spring and damper, forming a constrained three-body problem. The generalized coordinates are $q=(x_{1},y_{1},\theta_{1},x_{2},y_{2},\theta_{2},x_{3},y_{3},\theta_{3})$.

### v3

@@ 2.4 Qualitative Analysis of Error Propagation in Integration
$\Phi_{\Theta}(Z_{n}+E_{n})=\Phi_{\Theta}(Z_{n})+\frac{\partial\Phi_{\Theta}}{\partial Z}\bigg|_{Z_{n}}E_{n}+\mathcal{O}(\lVert E_{n}\rVert^{2})\;,$
where $\tfrac{\partial\Phi_{\Theta}}{\partial Z}\big|_{Z_{n}}$ denotes the Jacobian of the learned flow map $\Phi_{\Theta}$ with respect to the state, evaluated at $Z_{n}$. Consequently, one obtains
$E_{n+1}=\underbrace{\big(\Phi_{\Theta}(Z_{n})-\Phi(Z_{n})\big)}_{\text{local error}}+\underbrace{\frac{\partial\Phi_{\Theta}}{\partial Z}\bigg|_{Z_{n}}E_{n}}_{\text{propagated error}}+\mathcal{O}(\lVert E_{n}\rVert^{2})\;.$
Qualitatively, each step combines a local error (discrepancy between $\Phi_{\Theta}$ and $\Phi$) and a propagated error (amplification of the previous error by the solver Jacobian), showing how small inaccuracies in the learned acceleration field can accumulate over time, and thus underscoring the need for accurate acceleration targets during training.
To connect integration error with training error and differentiation-induced label error, we decompose the one-step defect. Let the true augmented dynamics be $\dot{Z}=F(Z)$ and let $\widehat{F}_{\Theta}(Z)$ denote the learned vector field (FNODE). Let $\Psi_{\Delta t}(\cdot;F)$ denote one step of a numerical integrator with step size $\Delta t$ applied to the vector field $F$. In this view, $\Phi(Z)=\Psi_{\Delta t}(Z;F)$ and $\Phi_{\Theta}(Z)=\Psi_{\Delta t}(Z;\widehat{F}_{\Theta})$ represent the corresponding one-step maps. The true and learned rollouts satisfy $Z_{n+1}=\Psi_{\Delta t}(Z_{n};F)$ and $\widehat{Z}_{n+1}=\Psi_{\Delta t}(\widehat{Z}_{n};\widehat{F}_{\Theta})$, and the global error $E_{n}=\widehat{Z}_{n}-Z_{n}$ evolves as $E_{n+1}=\Psi_{\Delta t}(Z_{n}+E_{n};\widehat{F}_{\Theta})-\Psi_{\Delta t}(Z_{n};F)$.
Adding and subtracting $\Psi_{\Delta t}(Z_{n};\widehat{F}_{\Theta})$ yields $E_{n+1}=\delta_{n}+\bigl(\Psi_{\Delta t}(Z_{n}+E_{n};\widehat{F}_{\Theta})-\Psi_{\Delta t}(Z_{n};\widehat{F}_{\Theta})\bigr)$, where $\delta_{n}:=\Psi_{\Delta t}(Z_{n};\widehat{F}_{\Theta})-\Psi_{\Delta t}(Z_{n};F)$ is the one-step defect at the true state. For an order-$p$ integrator, the defect admits a bound of the form $\|\delta_{n}\|\leq C_{\mathrm{int}}\Delta t^{p+1}+C_{\mathrm{fld}}\Delta t\,\|\widehat{F}_{\Theta}(Z_{n})-F(Z_{n})\|$, i.e., it combines the local truncation error of the time integrator and the effect of vector-field mismatch over one step.
The learned-field error has two contributions. FNODE is trained on numerically differentiated acceleration labels $\tilde{a}=a+e_{\mathrm{diff}}$, so $\widehat{F}_{\Theta}-F=(\widehat{F}_{\Theta}-\tilde{F})+(\tilde{F}-F)$, where $\tilde{F}$ is the vector field consistent with the differentiated labels. The first term is controlled by the training objective; in particular, if $L(\Theta)=\tfrac{1}{N}\sum_{i}\|\widehat{F}_{\Theta}(Z_{i})-\tilde{F}(Z_{i})\|_{2}^{2}$, then $\varepsilon_{\mathrm{train}}:=\sqrt{L(\Theta)}$ provides an RMS proxy for the learned-field mismatch on the training distribution. The second term reflects differentiation artifacts and can be assessed through the diagnostics introduced in Section 2.3.2, such as the stencil-disagreement indicator $\eta^{\prime}$ and the trajectory reconstruction residual $\eta_{z}$.
Finally, the second term in $E_{n+1}$ captures error propagation through integration. If the integrator is stable and satisfies $\bigl\|\tfrac{\partial}{\partial Z}\Psi_{\Delta t}(Z;\widehat{F}_{\Theta})\bigr\|\leq 1+C\Delta t$ over the rollout region, then $\|\Psi_{\Delta t}(Z_{n}+E_{n};\widehat{F}_{\Theta})-\Psi_{\Delta t}(Z_{n};\widehat{F}_{\Theta})\|\leq(1+C\Delta t)\|E_{n}\|$, which implies the recursion $\|E_{n+1}\|\leq(1+C\Delta t)\|E_{n}\|+\|\delta_{n}\|$. Consequently, $\|E_{n}\|$ grows at most like $(1+C\Delta t)^{n}$, and for $T=n\Delta t$ one obtains $\|E(T)\|\lesssim\exp(CT)\bigl(\|E(0)\|+\sum_{k}\|\delta_{k}\|\bigr)$. This shows explicitly how integration error accumulation couples to (i) the integrator order and step size through $C_{\mathrm{int}}\Delta t^{p+1}$, (ii) the training error through $\varepsilon_{\mathrm{train}}$, and (iii) differentiation-induced label error through the diagnostics $\eta^{\prime}$ and $\eta_{z}$.
@@ 2.5 Loss Function and Optimization
The FNODE loss directly supervises accelerations. Given predicted accelerations $\hat{\ddot{\mathbf{z}}}_{i}=f(\mathbf{z}_{i},\dot{\mathbf{z}}_{i};\mathbf{\Theta})$ and reference accelerations $\ddot{\mathbf{z}}_{i}$ obtained from trajectory data via FFT- or FD-based differentiation, we minimize the mean squared error (MSE)
$\mathcal{L}(\mathbf{\Theta})=\frac{1}{N}\sum_{i=0}^{N-1}\lVert\hat{\ddot{\mathbf{z}}}_{i}-\ddot{\mathbf{z}}_{i}\rVert^{2}_{2}=\frac{1}{N}\sum_{i=0}^{N-1}\lVert f(\mathbf{z}_{i},\dot{\mathbf{z}}_{i};\mathbf{\Theta})-\ddot{\mathbf{z}}_{i}\rVert^{2}_{2}.$ (24)
@@
$\mathbf{\Theta}^{*}=\arg\min_{\mathbf{\Theta}}\mathcal{L}(\mathbf{\Theta}).$ (25)
We train FNODE using the Adam optimizer [51]. The training algorithm is shown in A. A discussion of the training costs is deferred to Section 4.
@@ 3 Numerical Experiments
3.1 Single-Mass-Spring System
An undamped single mass–spring system serves as an energy-conserving benchmark for long-horizon prediction. Its dynamics are governed by
$\ddot{x}=-\tfrac{k}{m}x\;,$ (26)
where $x$ denotes the displacement from equilibrium, $m=10$ kg, and $k=50$ N/m. The system is initialized with $x(0)=1$ m and $\dot{x}(0)=0$, and simulated with step size $\Delta t=0.01$ s for 3000 steps (30 s). The first 300 steps are used for training and the remaining 2700 steps are reserved for extrapolation.
To highlight the system’s energy-conserving structure, we also express the dynamics in Hamiltonian form. Let $q=x$ be the generalized position and $p=m\dot{q}$ be the generalized momentum. The total energy (Hamiltonian) is
$T(p)=\frac{p^{2}}{2m},\qquad V(q)=\tfrac{1}{2}kq^{2},\qquad H(p,q)=T(p)+V(q)\;,$ (27)
which induces Hamilton’s equations
$\dot{q}=\frac{\partial H}{\partial p}=\frac{p}{m},\qquad\dot{p}=-\frac{\partial H}{\partial q}=-kq\;.$ (28)
We compare FNODE against structure-preserving baselines, including the Hamiltonian Neural Network (HNN) and Lagrangian Neural Network (LNN) [37]. HNN learns a scalar Hamiltonian $H_{\theta}(p,q)$ and obtains time derivatives through Hamilton’s equations, while LNN learns a Lagrangian $L_{\theta}(q,\dot{q})=T(\dot{q})-V(q)$ and predicts accelerations via the Euler–Lagrange equation. We also consider the Runge-Kutta 4th order (RK4) integrator as a baseline, which infers the future steps through the exact dynamics from Eq. (26) without learning.
Figure 2: Phase-space trajectories and energy evolution for the single mass–spring system. The left panel shows the phase portrait ($v$ vs. $x$), while the right panel shows the total energy. The energy drift ratios are: FNODE_RK4 ($15.70\%$), FNODE_LF ($0.46\%$), RK4 ($13.39\%$), HNN ($0.98\%$), and LNN ($0.24\%$). The mse for each model is provided: FNODE_RK4 ($\epsilon=7.50\text{e-}2$), FNODE_LF ($\epsilon=9.66\text{e-}5$), RK4 ($\epsilon=2.02\text{e-}3$), HNN ($\epsilon=5.51\text{e-}2$), and LNN ($\epsilon=2.27\text{e-}4$). The ground truth is provided by the analytical solution of the system.
Table 2: Hyper-parameters for the single mass spring system.
Hyper-parameters Model
FNODE HNN LNN
No. of hidden layers 2 2 2
No. of nodes per hidden layer 256 256 256
Max. epochs 450 30000 400
Initial learning rate 1e-3 1e-3 1e-3
Learning rate decay 0.98 0.98 0.98
Activation function tanh tanh tanh
Loss function MSE MSE MSE
Optimizer Adam Adam Adam
Figure 2 compares long-term rollout behavior over the 30-second trajectory. In phase space, all methods remain close to the ground truth closed orbit, indicating that they capture the periodic motion. However, the energy trajectories highlight marked differences in stability. Specifically, structure-aware modeling and numerical integration, e.g., use of the Hamiltonian or Lagrangian structure or using a symplectic integrator (leap-frog in our implementation – FNODE_LF), mitigate energy drift, see Fig. 2. HNN and LNN maintain near-constant energy throughout the rollout. In contrast, the RK4 integrator shows tangible energy drift, both when used directly on the equations of motion or in conjunction with the FNODE model (see FNODE_RK4).
@@ 3.2 Single-Mass-Spring-Damper System
The equation of motion for the single-mass-spring-damper system is given by
@@
where $m=10$ kg, $k=50$ N/m, and $d=2$ Ns/m. The system is initialized with $x(0)=1$ m, $\dot{x}(0)=0$, and solved by a Runge-Kutta 4th order (RK4) formula at $\Delta t=0.01$ s. A trajectory of 300 steps was used for training, with an additional 100 steps reserved for extrapolation. The model hyperparameters are listed in Table 3.
@@
Figure 3 compares predicted displacement and velocity. All methods fit well within the training interval, but differences emerge in extrapolation. FNODE maintains close agreement with the ground truth while MBD-NODE degrades more quickly. LSTM fails to capture the dissipative behavior, and FCNN diverges entirely. Phase-space trajectories in Figure 4 confirm FNODE’s stable long-term prediction, in contrast to the drift or stagnation observed in the baselines.
Figure 3: Temporal evolution of state variables for the single-mass-spring-damper system. ID stands for “in distribution” and OOD stands for “out of distribution”. The left and right columns depict the displacement ($x$) and velocity ($\dot{x}$), respectively. Dashed lines indicate model predictions over the training interval ($t\in[0,3]$), while dotted lines represent extrapolated predictions on the test interval. The mean squared error (MSE) for each model is provided: FNODE ($\epsilon=7.6\text{e-}5$), MBD-NODE ($\epsilon=1.4\text{e-}3$), LSTM ($\epsilon=1.5\text{e-}2$), and FCNN ($\epsilon=1.6\text{e-}1$). The ground truth is provided by the analytical solution of the system.
Figure 4: Comparative phase-space trajectories ($\dot{x}$ vs. $x$) for the single-mass-spring-damper system. Model performance on the training and test data is differentiated by dashed and dotted lines, respectively. The subplots correspond to the predictions of (a) FNODE, (b) MBD-NODE, (c) LSTM, and (d) FCNN.
@@ 3.3 Multiscale Triple-Mass-Spring-Damper System
We next consider a three-mass-spring-damper system with strong scale separation (Figure 5). The masses are $m_{1}=100$ kg, $m_{2}=10$ kg, and $m_{3}=1$ kg. All springs have stiffness $k_{i}=50$ N/m and all dampers $d_{i}=2$ Ns/m. The governing equations are
$\begin{aligned} \ddot{x}_{1}&=-\tfrac{k_{1}}{m_{1}}x_{1}-\tfrac{d_{1}}{m_{1}}\dot{x}_{1}+\tfrac{k_{2}}{m_{1}}(x_{2}-x_{1})+\tfrac{d_{2}}{m_{1}}(\dot{x}_{2}-\dot{x}_{1})\\ \ddot{x}_{2}&=-\tfrac{k_{2}}{m_{2}}(x_{2}-x_{1})-\tfrac{d_{2}}{m_{2}}(\dot{x}_{2}-\dot{x}_{1})+\tfrac{k_{3}}{m_{2}}(x_{3}-x_{2})+\tfrac{d_{3}}{m_{2}}(\dot{x}_{3}-\dot{x}_{2})\\ \ddot{x}_{3}&=-\tfrac{k_{3}}{m_{3}}(x_{3}-x_{2})-\tfrac{d_{3}}{m_{3}}(\dot{x}_{3}-\dot{x}_{2})\end{aligned}\;.$ (30)
Figure 5: Triple mass-spring-damper system. The setup is similar to the single mass-spring-damper system, except for the addition of two more masses, springs, and dampers. The system admits an analytical solution that was used for accuracy studies.
The system was simulated with RK4 at a fixed step of $\Delta t=0.01$ s. Training trajectories consisted of 300 steps from the initial state $x_{1}=1$, $x_{2}=2$, $x_{3}=3$, $\dot{x}_{1}=\dot{x}_{2}=\dot{x}_{3}=0$ (SI units). Performance was then evaluated by extrapolation for an additional 100 steps. Model hyperparameters are listed in Table 4.
@@
Figure 6 shows the trajectories of $x_{1},x_{2},x_{3}$ and $\dot{x}_{1},\dot{x}_{2},\dot{x}_{3}$ during training and testing. All models fit the training data, though FCNN exhibits minor spurious oscillations. In extrapolation ($t>3$), FNODE achieves the best agreement with ground truth ($\epsilon=7.4\text{e-}4$). MBD-NODE remains reasonable but drifts due to solver error accumulation. LSTM largely repeats training patterns and fails to generalize, while FCNN diverges under multiscale dynamics.
Figure 7 shows phase-space trajectories on the test set. FNODE tracks the dynamics well – the small discrepancies for the lighter masses are consistent with increased sensitivity to local inaccuracies in the learned acceleration field and their subsequent accumulation during rollout, as discussed in Section 2.4. MBD-NODE performs poorly for Mass 1, where low variance in the signal reduces its contribution to the loss. LSTM largely repeats training patterns and fails to generalize, while FCNN is inaccurate both in-distribution and out-of-distribution.
Figure 6: Predicted temporal evolution for the triple-mass-spring-damper benchmark. The left column presents the displacement ($x$) and the right column presents the velocity ($\dot{x}$) for each of the three masses. Dashed and dotted lines denote predictions on the training ($t\in[0,3]$) and test sets, respectively. The associated MSEs are: FNODE ($\epsilon=6.1\text{e-}3$), MBD-NODE ($\epsilon=5.4\text{e-}2$), LSTM ($\epsilon=2.5\text{e-}1$), and FCNN ($\epsilon=8.9\text{e-}2$).
@@ 3.4 Double Pendulum
We next evaluate FNODE on a classic example of a chaotic system – the double pendulum. The system is described by the Hamiltonian formulation (see B for details). We used $L_{1}=L_{2}=1$ m, $m_{1}=m_{2}=1$ kg, and $g=9.81$ m/s2. The initial state was $\theta_{1}=3\pi/7$, $\theta_{2}=3\pi/4$, and $\dot{\theta}_{1}=\dot{\theta}_{2}=0$. Trajectories were generated using RK4 with $\Delta t=0.01$ s for 300 training steps and 100 extrapolation steps. The model hyperparameters are listed in Table 5.
@@
Max. epochs 450 300 300 450
Initial learning rate 3e-3 1e-3 1e-3 3e-3
Learning rate decay 0.98 0.98 0.98 0.98
@@
Figure 9 shows the double pendulum trajectories. All models fit the training data, but diverge in extrapolation. FNODE, despite error accumulation in the ODE solver, achieves the best accuracy ($\epsilon=1.4\text{e-}1$).
Phase-space plots in Figure 10 confirm this trend: FNODE reproduces the main chaotic patterns, with only minor discrepancies. MBD-NODE loses accuracy in testing, LSTM largely repeats training patterns without generalization, and FCNN fails to capture the dynamics outside the training regime.
Figure 9: Dynamic response predictions for the double pendulum system. The left and right columns show the temporal evolution of angular position ($\theta$) and angular velocity ($\dot{\theta}$), respectively. Dashed lines represent the fit to the training data ($t\in[0,3]$), while dotted lines show the extrapolation on test data. The reported MSE values are: FNODE ($\epsilon=1.4\text{e-}1$), MBD-NODE ($\epsilon=2.3\text{e-}1$), LSTM ($\epsilon=7.6\text{e-}1$), and FCNN ($\epsilon=1.9\text{e}0$).
@@
For the double pendulum, which exhibits chaotic behavior in the regime considered here, “long-term” pointwise trajectory accuracy must be interpreted relative to the system’s intrinsic predictability horizon. In chaotic dynamics, small initial-condition or modeling errors often grow approximately exponentially (on average) at a rate set by the largest Lyapunov exponent $\lambda_{\max}$, motivating the Lyapunov time scale $T_{L}:=\tfrac{1}{\lambda_{\max}}$. Beyond horizons on the order of a few $T_{L}$, pointwise trajectory comparisons become increasingly stringent because even very small discrepancies are rapidly amplified by the dynamics. The rollout horizons reported in this section should therefore be interpreted as finite-horizon predictions (short to intermediate horizons) rather than claims of sustained pointwise accuracy arbitrarily far into the chaotic regime. Estimating $\lambda_{\max}$ for the specific operating regime (and reporting horizons in normalized units $t/T_{L}$) is an important direction for future work. Finally, even when long-horizon prediction is inherently challenging, FNODE (and related learned dynamics models) can still be useful in downstream tasks that rely on short-horizon predictive accuracy, e.g., when the prediction horizon of interest is short relative to $T_{L}$.
3.5 Slider Crank with Coulomb friction
Figure 11: Slider-crank mechanism with motor torque $\tau_{1}$, rotational dampers at all joints, and slider friction. The friction force is shown in red, see C, Eq. (52c), for its expression. It acts between body 3 and ground.
Next, FNODE’s performance is evaluated on a constrained slider-crank mechanism (Figure 11) under both frictionless and frictional conditions. The mechanism consists of three rigid bodies with a slider connected to a spring and damper, forming a constrained three-body problem. The generalized coordinates are $q=(x_{1},y_{1},\theta_{1},x_{2},y_{2},\theta_{2},x_{3},y_{3},\theta_{3})$; see C for details. The mechanism is defined by the following three bodies: